OpenStudy (clamin):

Please Help!! Medal! If tanX=3/4 , what is sin X??

OpenStudy (usukidoll):

you need the pythagorean theorem a^2+b^2 =c^2 so for tan x = 3/4 that's tan x = opposite / adjacent

OpenStudy (usukidoll):

so we need to solve for c 3^2+4^2 =c^2 9+16 =c^2 25=c^2 5 = c normally the answer would be 5 and -5 , but we can't have negative signs at all for geometric shapes.

OpenStudy (usukidoll):

Imma eat lunch

OpenStudy (wolf1728):

UsukiDoll The angle 'x' would be located in the lower left corner of the triangle.

OpenStudy (wolf1728):

sine (x) = opposite / hypotenuse sine (x) = 3 / c

OpenStudy (clamin):

So that would be 4/3??

OpenStudy (wolf1728):

No sine (x) = 3 / c What is 'c'?

OpenStudy (usukidoll):

it's asking for the sin x version we are given that tan x = 3/4 tan x is usually opposite/adjacent cos x is adjacent/hypotenuse sin x is opposite / hypotenuse your hypotenuse value is missing!

OpenStudy (usukidoll):

if the opposite -> 3 adjacent -> 4 our sin x is 3/ ? that ? is our unknown and we need the Pythagorean theorem for a^2+b^2=c^2

OpenStudy (usukidoll):



OpenStudy (usukidoll):

so using the Pythagorean theorem a^2+b^2=c^2 let your a = 3 and b = 4 3^2+(4)^2=c^2 9+16=c^2 25=c^2 5 = c that's your hypotenuse! since Hypotenuse is 5 and makes sense since it's supposed to be the longest side of the triangle Your sin x is opposite/ hypotenuse which is 3/5

OpenStudy (usukidoll):

sorry if I sounded a bit arrogant... but.. the reason why it's not 4/3 is because that's the cotangent so... sin x = opposite/hypotenuse cos x = adjacent/hypotenuse tan x = opposite/adjacent csc x, sec x, and cot x are backwards versions of what I've typed above tan x = opposite/adjacent (3/4) cot x = adjacent/opposite (4/3)